The market demand curve of a local pizza is QD= 100 − 4P. The total cost curve of Pat’s Pizza Kitchen is TC = 0.5Q2 + Q+5. Assuming Pat’s Pizza is doing business in a competitive industry and the price of the pizza is $10 for all firms. Using Excel to calculate the firm’s total revenue, total cost, and profit for q = 1 to q = 25 in increments of 1. (Note: your answers should be rounding decimals to the nearest whole number. Example: 12.3 => 12, 15.7=> 16):

At the price $10, how many units of output will the entire industry produce? ________ How many units of output will Pat’s Pizza produce to maximize profits (or minimize loss) in the short run? ________

Now, the industry turns into a monopoly.

Assuming the marginal cost curve stays the same as before. What is the profit-maximizing level of output? _________ What is total profit for the monopolist? ________ What is the price of the good the monopolist charges? ______ .
Answer #1
output (q) Price Total Revenue Total Cost Profit
1 10 10 6.5 3.5
2 10 20 9 11
3 10 30 12.5 17.5
4 10 40 17 23
5 10 50 22.5 27.5
6 10 60 29 31
7 10 70 36.5 33.5
8 10 80 45 35
9 10 90 54.5 35.5
10 10 100 65 35
11 10 110 76.5 33.5
12 10 120 89 31
13 10 130 102.5 27.5
14 10 140 117 23
15 10 150 132.5 17.5
16 10 160 149 11
17 10 170 166.5 3.5
18 10 180 185 -5
19 10 190 204.5 -14.5
20 10 200 225 -25
21 10 210 246.5 -36.5
22 10 220 269 -49
23 10 230 292.5 -62.5
24 10 240 317 -77
25 10 250 342.5 -92.5

QD = 100 - 4P

QD = 100 - 4(10)

QD = 100 - 40

QD = 60.

At the price $10, 60 units of output will the entire industry produce.

9 units of output will be produce to maximize the profit by Pat's Pizza.

Monopoly

TC = 0.5Q^2 + Q +5

MC = 0.5 (2)Q + 1.

MC = Q +1

Q = 100 - 4P
4P = 100 - Q
P = (100 -Q)/4
P = 25 - 0.25Q
TR = P Q
TR= (25 - 0.25Q) Q
TR = 25Q - 0.25Q^2

MR = 25 - 0.5Q

At equilibrium point, MR = MC

25 - 0.5Q = Q + 1

25 - 1 = Q + 0.5Q

24 = 1.5Q

Q = 24 / 1.5

Q = 16

P = 25 - 0.25Q

P = 25 - 0.25 (16)

P = 25 - 4

P = 21

Profit = TR - TC

Profit = P*Q - (0.5Q^2 + Q + 5)

Profit = (21) (16) - (0.5 (16)^2 + 16 + 5)

Profit = 336 - 149.

Profit = 187

The profit maximizing level of output is 16.

The total profit of monopolist is 187.

The price charged by the monopolist is 21.